40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 126 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  11284   Mon May 11 18:14:52 2015 ranaUpdateIMCMC_F calibration

I saw that entry, but it doesn't state what the calibration is in units of Hz/counts. It just gives the final calibrated spectrum.

  11283   Mon May 11 15:15:12 2015 manasaUpdateGeneralRan ASS for arms

Arm powers had drifted to ~ 0.5 in transmission.

X and Y arms were locked and ASS'd to bring the arm transmission powers to ~1.

  11282   Mon May 11 14:08:19 2015 manasaUpdateCDSc1lsp and c1sup removed?

I just found out that c1lsp and c1sup models no more exist on the FE status medm screens. I am assuming some changes were done to the models as well.

Earlier today, I was looking at some of the old medm screens running on Donatella that did not reflect this modification. 

Did I miss any elogs about this or was this change not elogged??

Quote:

I found the c1lsp and c1sup models not running anymore on c1lsc (white blocks for status lights on medm).

To fix this, I ssh'd into c1lsc. c1lsc status did not show c1lsp and c1sup models running on it.

I tried the usual rtcds restart <model name> for both and that returned error "Cannot start/stop model 'c1XXX' on host c1lsc".

I also tried rtcds restart all on c1lsc, but that has NOT brought back the models alive.

Does anyone know how I can fix this??

c1sup runs some the suspension controls. So I am afraid that the drift and frequent unlocking of the arms we see might be related to this.

 

P.S. We might also want to add the FE status channels to the summary pages.

 

  11281   Mon May 11 13:26:02 2015 manasaUpdateIMCMC_F calibration

The last MC_F calibration was done by Ayaka : Elog 7823

Quote:

And does anyone know what the MC_F calibration is?

 

  11280   Mon May 11 13:21:25 2015 manasaUpdateCDSc1lsp and c1sup not running

I found the c1lsp and c1sup models not running anymore on c1lsc (white blocks for status lights on medm).

To fix this, I ssh'd into c1lsc. c1lsc status did not show c1lsp and c1sup models running on it.

I tried the usual rtcds restart <model name> for both and that returned error "Cannot start/stop model 'c1XXX' on host c1lsc".

I also tried rtcds restart all on c1lsc, but that has NOT brought back the models alive.

Does anyone know how I can fix this??

c1sup runs some the suspension controls. So I am afraid that the drift and frequent unlocking of the arms we see might be related to this.

 

P.S. We might also want to add the FE status channels to the summary pages.

  11279   Mon May 11 12:17:19 2015 max isiHowToGeneralSummary pages

I have created a wiki page with introductory info about the summary page configuration: https://wiki-40m.ligo.caltech.edu/Daily summary help

We can also use that to collect tips for editing the configuration files, etc.

Quote:

I have set up new summary pages for the 40m: http://www.ligo.caltech.edu/~misi/summary/
This website shows plots (time series, spectra, spectrograms, Rayleigh statistics) of relevant channels and is updated with new data every 30 min.

The content and structure of the pages is determined by configuration files stored in nodus:/users/public_html/gwsumm-ini/ . The code looks at all files in that directory matching c1*.ini. You can look at the c1hoft.ini file to see how this works. Besides, a quick guide to the format can be found here http://www.ligo.caltech.edu/~misi/iniguide.pdf

Please look at the pages and edit the config files to make them useful to you. The files are under version control, so don’t worry about breaking anything.

Do let me know if you have any questions (or leave a comment in the pages).

 

  11278   Mon May 11 01:28:33 2015 ranaHowToComputer Scripts / Programssummary page URL change

Like Steve pointed out, the summary pages show that the y-arm transmission drifts a lot when locked. The OL summary page shows that this is all due to ITMY yaw.

Could be either that they coil driver / DAC is bad or that the suspension is poorly built. We need to dig into ITMY OL trends over long term to see if this is new or now.

Also, weather station needs a reboot. And does anyone know what the MC_F calibration is?

  11277   Sun May 10 13:54:41 2015 ranaHowToComputer Scripts / Programssummary page URL change

Also, EQ gave us a better (and not pwd protected) URL for the summary pages. Please replace your previous links with this new one:

https://nodus.ligo.caltech.edu:30889/detcharsummary/

  11276   Fri May 8 14:30:09 2015 SteveUpdateVACCC1 cold cathode gauges are baked now

CC1s  are not reading any longer. It is an attempt to clean them over the weekend at 85C

These brand new gauges "10421002" sn 11823-vertical, sn 11837 horizontal replaced 11 years old 421 http://nsei.missouri.edu/manuals/hps-mks/421%20Cold%20Cathode%20Ionization%20Guage.pdf  on 09-06-2012 http://nodus.ligo.caltech.edu:8080/40m/7441

Quote:

 

We have two cold cathode gauges at the pump spool and one  signal cable to controller. CC1  in horizontal position and CC1 in vertical position.  

CC1 h started not reading so I moved cable over to CC1 v

 

Attachment 1: cc1bake95C.jpg
cc1bake95C.jpg
  11275   Fri May 8 08:16:46 2015 SteveUpdateLSCdrift in Y arm

Why is the Y arm drifting so much?

The " PSL FSS Slow Actuator Adjust " was brought back to range from 1.5 to 0.3 yesterday as ususual. Nothing else was touched.

I'm not sure if the timing scale is working correctly on theses summery plots. What is the definition of today?

The y-arm became much better as I noticed it at 5pm

 

Attachment 1: 1goodday.png
1goodday.png
Attachment 2: Yishappy.png
Yishappy.png
  11274   Tue May 5 16:02:57 2015 SteveUpdateVACVac Summery Channels with discription

As it was requested by the Bos.

It would be nice to read from the epic screen C1:Vac-state_mon.......Current State: Vacuum Normal, valve configuration

Quote:

       Channel        

                    Function                                              Description                                             Interlock         
             

C1:Vac-P1_pressure   

 Main volume of 40m interferro meter        P=Pirani gauge, Pressure range: ATM-760  to 1e-4 Torr at 3 mT close V1 and PSL shutter
C1:Vac-P2_pressure  Maglev foreline pressure

Maglev is the main pump of our vacuum system below 500 mTorr

It's long term pressure has to be <500 mTorr                  

 at 6 Torr close V1 
C1:Vac-P3_pressure  annuloses

 Each chamber has it's own annulos. These small volumes are indipendent from main volume.     Their  pressure  ranges are <5 mTorr at vac. normal valve configuration.

                                          
C1:Vac-CC1_pressure  IFO main volume

CC1=cold cathode gauge (low emmision), Pressure range: 1e-4 to 1e-10 Torr,

In vac- normal configuration CC1= 2e-6 Torr

at 1e-5 Torr close VM1
C1:Vac-CC4_pressure  RGA  pressure In vac-normal configuration CC1=CC4  
 C1:Vac-N2pres  valve's drive pneumatic    

The N2 supply is regulated to 60-80 PSI out put at the auto cylinder changer.

 at 55 PSI close V1, at 45 PSI close all 
 It  does not exist yet 2 N2 cylinder sum pressure

Each cylinder pressure will be measured before the regulator and summed for warning message to be send

at 1000 PSI

 

  11273   Tue May 5 10:40:05 2015 ericqHowToComputer Scripts / ProgramsHow to get a web page running on Nodus

How to get your own web page running on Nodus

  1. On any martian machine, put your stuff in /users/public_html/$MYPAGE/
  2. On Nodus, run: ln -s /users/public_html/$MYPAGE /export/home/
  3. Your site is now available at https://nodus.ligo.caltech.edu:30889/$MYPAGE/
  4. If you want to allow straight up directory listing to the entire internet, on Nodus run: sudoedit /etc/sites-available/nodus, and add the following lines towards the bottom:
<Directory /export/home/$MYPAGE>
    Options +Indexes
</Directory>
  11272   Mon May 4 12:42:34 2015 manasaSummaryGeneralDelay line frequency discriminator for FOL error signal

Koji suggested that I make a cosine fit for the curve instead of a linear fit.

I fit the data to V(f) = A + B cos(2\pi f_{b}L/v) 
where L - cable length asymmetry (27 cm) , fb - beat frequency and v - velocity of light in the cable (2*10m/s)

The plot with the cosine fit is attached. 

Fit coefficients (with 95% confidence bounds):
       A =      0.4177  (0.3763, 0.4591)
       B =       2.941  (2.89, 2.992)

Attachment 1: DFD_cosfit.png
DFD_cosfit.png
  11271   Mon May 4 12:35:49 2015 ranaUpdateLSCdrift in Y arm

http://www.ligo.caltech.edu/~misi/summary/day/20150504/

I left the arms locked last night. Looks like the drift in the Y arm power is related to the Y arm control signal being much bigger than X.

Why would it be that Y > X  ?

  11270   Mon May 4 10:21:09 2015 manasaSummaryGeneralDelay line frequency discriminator for FOL error signal

Attached is the schematic of the analog DFD and the plot showing the zero-crossing for a delay line length of 27cm. The bandwidth for the linear output signal obtained roughly matches what is expected from the length difference (370MHz) .

We could use a smaller cable to further increase our bandwidth. I propose we use this analog DFD to determine the range at which the frequency counter needs to be set and then use the frequency counter readout as the error signal for FOL.

 

Attachment 1: DFD.png
DFD.png
Attachment 2: DFD_resp.png
DFD_resp.png
  11269   Sun May 3 19:40:51 2015 ranaUpdateASCSunday maintenance: alignment, OL center, seismo, temp sensors

X arm was far out in yaw, so I reran the ASS for Y and then X. Ran OK; the offload from ASS outputs to SUS bias is still pretty violent - needs smoother ramping.

After this I recentered the ITMX OL- it was off by 50 microradians in pitch. Just like the BS/PRM OLs, this one has a few badly assembled & flimsly mounts. Steve, please prepare for replacing the ITMX OL mirror mounts with the proper base/post/Polaris combo. I think we need ~3 of them. Pit/yaw loop measurements attached.

Based on the PEM-SEIS summary page, it looked like GUR1 was oscillating (and thereby saturating and suppressing the Z channel). So I power cycled both Guralps by turning off the interface box for ~30 seconds and the powering back on. Still not fixed; looks like the oscillations at 110 and 520 Hz have moved but GUR2_X/Y are suppressed above 1 Hz, and GUR1_Z is suppressed below 1 Hz. We need Jenne or Zach to come and use the Gur Paddle on these things to make them OK.

From the SUS-WatchDog summary page, it looked like the PRM tripped during the little 3.8 EQ at 4AM, so I un-tripped it.

Caryn's temperature sensors look like they're still plugged in. Does anyone know where they're connected?

Attachment 1: itmx_ol_loops_150503.png
itmx_ol_loops_150503.png
Attachment 2: Gur_150503.png
Gur_150503.png
  11268   Sun May 3 01:04:19 2015 ranaSummaryPEMSeismo signals are bad

https://ldas-jobs.ligo.caltech.edu/~max.isi/summary/day/20150502/pem/seismic/

Looks like some of our seismometers are oscillating, not mounted well, or something like that. No reason for them to be so different.

Which Guralp is where? And where are our accelerometers mounted?

  11267   Fri May 1 20:33:31 2015 ranaSummaryComputer Scripts / Programsproblems with new restart procedures for elogd and apache

Same thing again todaysad. So I renamed the /etc/init/elog.conf so that it doesn't keep respawning bootlessly. Until then restart elog using the start script in /cvs/cds/caltech/elog/ as usual.

I'll let EQ debug when he gets back - probably we need to pause the elog respawn so that it waits until nodus is up for a few minutes before starting.

Quote:

Since the nodus upgrade, Eric/Diego changed the old csh restart procedures to be more UNIX standard. The instructions are in the wiki.

After doing some software updates on nodus today, apache and elogd didn't come back OK. Maybe because of some race condition, elog tried to start but didn't get apache. Apache couldn't start because it found that someone was already binding the ELOGD port. So I killed ELOGD several times (because it kept trying to respawn). Once it stopped trying to come back I could restart Apache using the Wiki instructions. But the instructions didn't work for ELOGD, so I had to restart that using the usual .csh script way that we used to use.

 

  11266   Fri May 1 16:42:42 2015 ranaUpdateDAQPEM Slow channels added to saved frames

Still processing, but I think it should work fine once we have a day of data. Until then, here's the summary pages so far, including Vac channels:

http://www.ligo.caltech.edu/~misi/summary/day/20150501/pem/

  11265   Fri May 1 13:22:08 2015 ericqUpdateDAQPEM Slow channels added to saved frames

Rana asked me to include add slow outputs (OUT16) of the seismometer BLRMS channels to the frames. 

All of the PEM slow channels are already set up in c1/chans/daq/C1EDCU_PEM.ini, but up to this point, daqd had no knowledge of this file, since it wasn't included in c1/target/fb/master, which defines all the places to look for files describing channels to be written to disk. This file already includes lines for C1EDCU_LSC.ini and such, which from old elogs, looks like was set up by hand for subsystems we care about. 

Hence, since we now care about slow trends for the PEM subsystem, I have added a line to the daqd master file to tell it to save the PEM slow channels. This looks to have increased the size of the individual 16 second frame files from 57MB to 59MB, which isn't so bad.

  11264   Thu Apr 30 16:30:25 2015 SteveUpdateVACN2 pneumatic pressure watch set up

We have 2 transduser PX303-3KG5V   http://www.omega.com/pressure/pdf/PX303.pdf They  will be installed on the out put of the N2 cylinders to read the supply pressure.

I will order one DC power supply  http://www.omega.com/pptst/PSU93_FPW15.html      PSU-93

One full cylinder pressure is ~ 2400 PSI max so two of them will give us ~9Vdc

The email reminder should be send at 1000 PSI  =  1.8 V

 

 

 

 

  11263   Wed Apr 29 18:12:42 2015 ranaUpdateComputer Scripts / Programsnodus update

Installed libmotif3 and libmotif4 on nodus so that we can run dataviewer on there.

Also, the lscsoft stuff wasn't installed for apt-get, so I did so following the instructions on the DASWG website:

https://www.lsc-group.phys.uwm.edu/daswg/download/repositories.html#debian

Then I installed libmetaio1, libfftw3-3. Now, rather than complain about missing librarries, diaggui just silently dies.

Then I noticed that the awggui error message tells us to use 'ssh -Y' instead of 'ssh -X'. Using that I could run DTT on nodus from my office.

  11262   Tue Apr 28 09:49:26 2015 SteveUpdateVACVac Summery Channels

 

 

       Channel        

                    Function                                          Interlock action           
            

C1:Vac-P1_pressure   

 IFO vac envelope pressure           at 3 mT close V1 and PSL shutter
C1:Vac-P2_pressure  Maglev foreline pressure                      at 6 Torr close V1 
C1:Vac-P3_pressure  annuloses    
C1:Vac-CC1_pressure  IFO pressure   at 1e-5 Torr close VM1
C1:Vac-CC4_pressure  RGA  pressure    
 C1:Vac-N2pres  valve's drive  pneumatic 60-80PSI    

 

at 55 PSI close V1, at 45 PSI close all 
 It  does not exist yet 2 N2 cylinder sum pressure  

 

  11261   Mon Apr 27 21:42:07 2015 ranaUpdateVACSummary pages

We want to have a VAC page in the summaries, so Steve - please put a list of important channel names for the vacuum system into the elog so that we can start monitoring for trouble.

Also, anyone that has any ideas can feel free to just add a comment to the summary pages DisQus comment section with the 40m shared account or make your own account.

  11260   Mon Apr 27 14:37:55 2015 SteveUpdateVACN2 pneumatic pressure watch

 

Quote:

Based on Jenne's chiara disk usage monitoring script, I made a script that checks the N2 pressure, which will send an email to myself, Jenne, Rana, Koji, and Steve, should the pressure fall below 60psi. I also updated the chiara disk checking script to work on the new Nodus setup. I tested the two, only emailing myself, and they appear to work as expected. 

The scripts are committed to the svn. Nodus' crontab now includes these two scripts, as well as the crontab backup script. (It occurs to me that the crontab backup script could be a little smarter, only backing it up if a change is made, but the archive is only a few MB, so it's probably not so important...)

This watch script gives little time to replace N2 cylinder. When the regulated supply drops below 60 psi the cylinder pressure is 60 psi too.

It is more of a statement that V1 is closed and act accordingly. It's only practical if you are in the lab.

Rana pointed it out correctly that we need this message 24 hrs before it happens. This requires monitoring  of total supplies , not the regulated one.

So we need pressure transducers on each nitrogen cylinder, before the regulator. The sum of the two N2 cylinder when they are full 4000 - 4500 psi

The first email should be send out at 1000 psi as sum of the two cylinders. This means that you have  1 day to replace nitrogen cylinder.

 Most of the time the daily consumption is 750 +-50 psi   

However sometimes  this variation goes up   ~750 +-150 psi

  11259   Mon Apr 27 09:09:15 2015 SteveUpdatePEMair cond filters checked

 

Quote:

 

Quote:

Yesterday morning was dusty. I wonder why?

The PRM sus damping was restored this morning.

Yesterday afternoon at 4 the dust count peaked 70,000 counts

Manasa's alergy was bad at the X-end yesterday. What is going on?

There was no wind and CES neighbors did not do anything.

Air cond filters checked by Chris. The 400 days plot show 3 bad peaks at 1-20, 2-5 & 2-19

Attachment 1: 400d.png
400d.png
  11258   Mon Apr 27 01:13:08 2015 JenneUpdateLSCPRCL angular FF not working, no locking :(

I'm sad.  And frustrated. 

The PRCL angular feed forward is not working, and without it I am having a very difficult time keeping the PRMI locked while the arms are at high power (either buzzing, or the one time I got stable high power partway through the transition).  Obviously if the PRMI unlocks once CARM and DARM are mostly relying on the REFL signals, I lose the whole IFO. 

Q and I had been noticing over the last few weeks that the angular feed forward wasn't seeming quite as awesome as it did when I first implemented it.  We speculated that this was likely because we had started DC coupling the ITM optical levers, which changes the way seismic motion is propagated to cavity axis motion (since the ITMs are reacting differently).

Anyhow, today it does not work at all.  It just pushes the PRM until the PRMI loses lock. I am worried that, even though Rana re-tuned the BS and PRM oplev servos to be very similar to how they used to be, there is enough of a difference (especially when compounded with the DC coupled ITMs) that the feed forward transfer functions just aren't valid anymore.

Since this prevents whole IFO locking, I spent some time trying to get it back under control, although it's still not working. 

I remeasured the actuator transfer function of how moving PRM affects the sideband spot at the QPD, in the PRMI-only situation.  I didn't make a comparison plot for the yaw degree of freedom, but you can see that the pitch transfer function is pretty different below ~20Hz, which is the whole region that we care about.  In the plot below, black is from January (PRMI-only, no DC-coupled ITMs) and blue is from today (PRMI-only, with DC-coupled ITMs, and somewhat different BS/PRM oplev setup):

Pitch_oldVsNew.pdf

I calculated new Wiener filters, and tried to put them in, but sometimes (and I don't understand what the pattern is yet) I get "error" in the Alternate box, rather than the zpk version of my sos filter.  It seems to go away if you use fewer and fewer poles for fitting the Wiener filters, but then the fit is so poor that you're not going to get any subtraction (according to the residual estimation plot that uses the fitted filters rather than the ideal Wiener filters). The pitch filters could only handle 6 poles, although the yaw filters were fine with 20.

The feed forward just keeps pushing the PRM away though.  I flipped the signs on the Wiener filters, I tried recalculating without the actuator pre-filtering, I don't know why it's failing.  But, I'm not able to lock the interferometer.  Which sucks, because I was hoping to finally get most of my noise coupling measurements done today.

 

Attachment 1: Pitch_oldVsNew.pdf
Pitch_oldVsNew.pdf
  11257   Sun Apr 26 20:10:10 2015 max isiHowToGeneralSummary pages

I have set up new summary pages for the 40m: http://www.ligo.caltech.edu/~misi/summary/
This website shows plots (time series, spectra, spectrograms, Rayleigh statistics) of relevant channels and is updated with new data every 30 min.

The content and structure of the pages is determined by configuration files stored in nodus:/users/public_html/gwsumm-ini/ . The code looks at all files in that directory matching c1*.ini. You can look at the c1hoft.ini file to see how this works. Besides, a quick guide to the format can be found here http://www.ligo.caltech.edu/~misi/iniguide.pdf

Please look at the pages and edit the config files to make them useful to you. The files are under version control, so don’t worry about breaking anything.

Do let me know if you have any questions (or leave a comment in the pages).

  11256   Sun Apr 26 15:34:34 2015 JenneUpdateSUSPRM oplev centered

After last week's work on the BS/PRM oplev table, I think the PRM oplev got centered while the PRM was misaligned.  With the PRM aligned, the oplev spot was not on the QPD.  It has been centered.

  11255   Sun Apr 26 15:05:35 2015 JenneUpdateASCunBroken Xass?

Thank you both.

I have updated the .snap file, so that it'll use these parameters, as Rana left them.  Also, so that the "unfreeze" script works without changes (since it wants to make the overall gain 1), I have changed the Xarm input matrix elements from 1 to 0.1, for all of them.  This should be equivalent to the overall gain being 0.1.

  11254   Sun Apr 26 14:17:40 2015 JenneUpdateLSCPOXDC, POYDC unplugged for now

I have unplugged POXDC and POYDC from their whitening inputs.  They have labels on them which whitening channel they belong to (POY=5, POX=6) on the DCPD whitening board.

TT3_LR's DAC output is Tee-ed, going to the POYDC input and also to an SR560 near the Marconi.

TT4_LR's DAC output is Tee-ed, going to the POXDC input and also to the CM board's ExcB input.

  11253   Sun Apr 26 01:10:18 2015 ranaUpdateASCunBroken Xass?

Today I tried some things, but basically, lowering the input gain by 10 made the thing stable. In the attached screenshotstrip, you can see what happens with the gain at 1. After a few cycles of oscillation, I turned the gain back to 0.1.

There still is an uncontrolled DoF, but I that's just the way it is since we only have one mirror (the BS) to steer into the x arm once the yarm pointing is fixed.

Along the way, I also changed the phase for POX, just in case that was an issue. I changed it from +86 to +101 deg. The attached spectra shows how that lowered the POX_Q noise.

I also changed the frequencies for ETM_P/Y dither from ~14/18 Hz to 11.31/14.13 Hz. This seemed to make no difference, but since the TR and PO signals were quieter there I left it like that.

This is probably OK for now and we can tune up the matrix by measuring some sensing matrix stuff again later.

Attachment 1: xasstune_150426.png
xasstune_150426.png
Attachment 2: xassnoise.pdf
xassnoise.pdf
  11252   Sun Apr 26 00:56:21 2015 ranaSummaryComputer Scripts / Programsproblems with new restart procedures for elogd and apache

Since the nodus upgrade, Eric/Diego changed the old csh restart procedures to be more UNIX standard. The instructions are in the wiki.

After doing some software updates on nodus today, apache and elogd didn't come back OK. Maybe because of some race condition, elog tried to start but didn't get apache. Apache couldn't start because it found that someone was already binding the ELOGD port. So I killed ELOGD several times (because it kept trying to respawn). Once it stopped trying to come back I could restart Apache using the Wiki instructions. But the instructions didn't work for ELOGD, so I had to restart that using the usual .csh script way that we used to use.

  11251   Sun Apr 26 00:08:56 2015 ranaHowToelogTroubles with putting plots in external locations

If it all possible, don't use links to your home directory. Its not robust. It would be like if you clicked on your Google Music and it told you to ask me to sing that song to you. Imagine that on auto-repeat next time your fancy-bone itches.

Attachment 1: 48.png
48.png
  11250   Sat Apr 25 22:17:49 2015 ranaUpdateCDSMXstream restart script working (beta)

Since python from crontab seemed intractableangry, I replaced autoMX.py with a soft link that points at autoMX.sh.

This is a simple BASH scriptcool that looks at the LSC FB stat (C1:DAQ-DC0_C1LSC_STATUS), and runs the restart mxstream script if its non-zero.

So far its run 5 times successfullylaugh. I guess this is good enough for now. Later on, someone ought to make it loop over other FE, but this ought to catch 99% of the FB issues.

  11249   Sat Apr 25 18:50:47 2015 ericqUpdateVACPressure watch script broken

Ugh, this turns out to be because cron doesn't source the controls bashrc that defines where to find caget and all that jazz that many commands depend on. This is probably also why the AutoMX cron job isn't working either. 

Also, cron automatically emails everything from stderr to the email address that is configured for the user, which is why the n2 script blew up the foteee account and why the AutoMX script was blowing up my email yesterday. This can be avoided by doing something like this in the crontab:

0 8 * * * /bin/somecommand >> somefile.log 2>&1

(The >> part means that the standard output is appended to some log file, while the 2>&1 means send the standard error stream to the same place as stdout)

I made this change for the n2 script, so the foteee email account should be safe from this script. I haven't figured out the right way to set up cron to have all the right $PATH and other environment stuff, such as epics may need, so the script is still not working. 

  11248   Sat Apr 25 03:32:45 2015 KojiUpdateASCBroken Xass?

I spent a day to fix the XARM ASS, but no real result. If the input of the 6th DOF servo is turned off, the other error signals are happy
to be squished to around their zeros. So this gives us some sort of alignment control. But obviously a particular combination of the
misalignment is left uncontrolled.

This 6th DOF uses BS to minimize the dither in ITMX yaw. I tired to use the other actuators but failed to have linear coupling between
the actuator and the sensor.


During the investigation, I compared TRX/TRY power spectra. TRX had a bump at 30Hz. Further investigation revealed that the POX/POY
had a big bump in the error signals. The POX/POY error signals between 10-100Hz were coherent. This means that this is coming from
the frequency noise stabilized with the MC. (Is this frequency noise level reasonable?)

The mysterious discovery was that the bump in the transmission exist only in TRX. How did the residual frequency noise cause
the intensity noise of the transmission? One way is the PDH offset.

Anyway, Rana pointed out that IMC WFS QPDs had large spot offsets. Rana went to the AS table and fixed the WFS spot centering.
This actually removed the bump in TRX although we still don't know the mechanism of this coupling.

The bump at 30Hz was removed. However, the ASS issue still remains.

  11247   Sat Apr 25 00:20:16 2015 ranaUpdateCDSmegatron python autoMC cron

Upgraded python on megatron. Added lines to the crontab to run autoMX.py. Edited crontab to have a PYTHONPATH so that it can run .py stuff.

But autoMX.py is still not working from inside of cron, just from command line.

  11246   Fri Apr 24 23:40:15 2015 ranaUpdateSUSPRM and BS oplev laser replaced

Recently, Steve replaced the HeNe which was sourcing the BS & PRM OL. After replacement, no one checked the beam sizes and we've been living with a mostly broken BS OL. The beam spot on the QPD was so tiny that we were seeing the 'beam is nearly the size of the segment gap' effect.

Today I removed 2 of the lenses which were in the beam path: one removed from the common PRM/BS path, and one removed from the PRM path. The beams on both the BS & PRM got bigger. The BS beam is bigger by a factor of 7. I've increased the loop gains by a factor of 6 and now the UGFs are ~6 Hz. The loop gains were much too high with the small beam spots that Steve had left there. I would prefer for the beams to be ~1.5-2x smaller than they are now, but its not terrible.

Many of the mounts on the table are low quality and not constructed stably. One of the PRM turning mirror mounts twisted all the way around when I tried to align it. This table needs some help this summer.

In the future: never try locking after an OL laser change. Always redo the telescope and alignment and check the servo shape before the OL job is done.

Also, I reduced the height of the RG3.3 in the OL loops from 30 to 18 dB. The BS OL loops were conditionally stable before and thats a no-no. It makes it oscillate if it saturates.

Attachment 1: BSOL.pdf
BSOL.pdf
  11245   Fri Apr 24 21:31:20 2015 KojiSummaryCDSautomatic mxstream resetting

We were too much annoyed by frequent stall of mxstream. We'll update the RCG when time comes (not too much future).

But for now, we need an automatic mxstream resetting.

I found there is such a script already.

/opt/rtcds/caltech/c1/scripts/cds/autoMX

So this script was registered to crontab on megatron.
It is invoked every 5 minutes.

# Auto MXstream reset when it fails
0,5,10,15,20,25,30,35,40,45,50,55 * * * * /opt/rtcds/caltech/c1/scripts/cds/autoMX >> /opt/rtcds/caltech/c1/scripts/cds/autoMX.log

  11244   Fri Apr 24 18:13:36 2015 ranaUpdateGeneraltorque for 1/4-20

For 1/4-20 bolts made of 18-8 Stainless Steel, the recommended torque varies from 65-100 inch-pounds, depending upon the application, the lubrication, how loose the bolt is, if there's a washer, etc.

For our case, where we are going into a tapped, ferromagnetic stainless table, its less clear, but it will certainly by in the 60-80 range. This is close to the 5-6 foot-lbs that I recommended on Wednesday.

I've ordered 3 torque wrenches with 1/4" drive so that we can have one at each end and one in the toolbox near MC2. We'll indicate the recommended torque on there so that we can tighten everything appropriately.

  11243   Fri Apr 24 17:30:32 2015 JenneUpdateVACPressure watch script broken
Quote:

I made a script that checks the N2 pressure, which will send an email to myself, Jenne, Rana, Koji, and Steve, should the pressure fall below 60psi.

The script checking the N2 pressure is not working.  I signed into the foteee account to look at some of the picasa photos, and there are thousands of emails (one every 10 minutes for the past month!) with error messages.  Q, can you please make it stop (having errors)?

The error looks like it's mad about a "caget" command.  I don't have time to investigate further though.

  11242   Fri Apr 24 01:16:30 2015 JenneUpdateASCBroken Xass?

I ran the "off" script for the Xarm ASS, followed by the "on" script, and now the Xarm ASS doesn't work.  Usually we just run the freeze/unfreeze, but I ran the off/on scripts one time. 

Koji, if you have some time tomorrow, can you please look at it?  I am sorry to ask, but it would be very helpful if I could keep working on other things while the ASS is taken care of.

Steve, can you please find a cable that goes from the LSC rack to the IOO rack (1Y2 to 1X2), or lay a new one?  It must be one single long cable, without barrels sticking it together.  This will help me actuate on the Marconi using the LSC rack's DAC. 

Thank you!!

  11241   Thu Apr 23 23:07:23 2015 DugoliniFrogsALARMlaptops warning

Please!

Don't put laptops on the ISC Tables!

  11240   Thu Apr 23 21:05:23 2015 ranaUpdateComputer Scripts / ProgramsCDSutils upgrade undone

Q: please update this Wiki page with the go-back procedure:

https://wiki-40m.ligo.caltech.edu/CDSutils_Upgrade_Procedure

  11239   Thu Apr 23 15:40:41 2015 SteveUpdateGeneraltorque for 1/4-20

 

Few 1/4 -20 socket cap head screw with washers were tested for optimum torque.

QJR 117E Snap On  torque wrench was used. I found that 40 lb in was enough.

Looked up recommended values on the web later:

Our Thorlab SS 1/4-20 screw kits are SS 18-8 as DRY 70 inch / lbs max, lubricated  60 inch / lbs max

These numbers will varie with washers, material it's going into and so on!

BLACK-OXIDE Alloy Steel Socket Head Cap Screws can go much higher value

Thread Size 1/4"-20
Length 1/4"
Thread Length Full
Additional Specifications Black-Oxide Alloy Steel
RoHS Compliant

The standard among high-strength fasteners, these screws are stronger than Grade 8 steel screws. They have a minimum tensile strength of 170,000 psi. and a minimum Rockwell hardness of C37. Length is measured from under the head.

Inch screws have a Class 3A thread fit. They meet ASTM A574.

Black Oxide—Screws have been heat-treated for hardness, which results in a dark surface color.

 
The information in this 3-D model is provided for reference only. Details
 
 

We still do not know that what torque values we get best performance : minimum jiggel and drift etc.

After looking at these numbers I raise my recommendation to 50 inch / lbs on a std aplication.

Rana is next to calibrate his feelings and declare the right number.

Than Koji....and so on

Once we a number, than I buy more torque wrenches to fit it.

Attachment 1: snapon.jpg
snapon.jpg
Attachment 2: QJR117E.jpg
QJR117E.jpg
  11238   Thu Apr 23 08:43:40 2015 SteveUpdateElectronicsEG&G InGaAs diodes in stock

RFpds box is moved from RF cabinet E4 to clean cabinet S15

Inventory updated at https://wiki-40m.ligo.caltech.edu/RF_Pd_Inventory

Large Area InGaAs PIN Photodiode -- C30642GH      6 pieces in stock

Product Details
in: Photodiodes

 

Large Area InGaAs PIN Photodiode -- C30642GH -- View Larger Image

Large Area InGaAs PIN Photodiode with a 2,0 mm active diameter chip in TO-5 package with flat glass window

Large area InGaAs PIN photodiode with useful diameter of 2,0 mm in a T0-5 package with a flat glass window. The C30642GH provides high quantum efficiency from 800nm to 1700nm. It features high responsivity, high shunt resistance, low dark current, low capacitance for fast response time and uniformity within 2% across the detector active area.

  11237   Wed Apr 22 17:04:11 2015 ranaUpdateElectronicsMC REFL PD back from the dead

Just randomly found this old entry from 3 years ago. We should never have installed a GAP 2000 - they are an inferior type of InGaAs diode. We should add to our list replacing these with a 2 mm EG&G diode.

How many 2 mm EG&G InGaAs diodes do we have Steve? Can you please find a good clean diode case so that we can store them in the optics cabinet on the south arm?

Quote:

 [Yuta, Manasa]

We replaced the dead photodiode on MC REFL PD with a new one (GAP 2000). We measured the frequency response of the PD and tuned the resonant frequency using inductor L5 (in the circuit diagram) to be 29.575MHz - over an average of 10 measurements.

 

  11236   Wed Apr 22 14:56:18 2015 manasaSummaryGeneralDelay line frequency discriminator for FOL error signal

[Koji, Manasa]

Since the bandwidth of the fiber PD is ~ 1GHz, we could design the frequency discriminator to have a wider bandwidth (~ 500MHz). The output from the frequency discriminator could then be used to define the range setting of the frequency counter for readout or may be even error signal to the PID loop.

A test run for the analog DFD with cable length difference of 27cm gave a linear output signal with zero-crossing at ~206MHz.

Detailed schematic of the setup and plot (voltage vs frequency) will be updated shortly.

  11235   Wed Apr 22 11:48:30 2015 manasaSummaryGeneralDelay line frequency discriminator for FOL error signal

Since the Frequency counters have not been a reliable error signal for FOL PID loop, we will put together an analog delay line frequency dicriminator as an alternative method to obtain the beat frequency.

The configuration will be similar to what was done in elog 4254 in the first place.

For a delay line frequency dicriminator, the output at the mixer is proportional to cos(\theta_{b}) where \theta_{b} = 2 \pi f_{b}L/v

L - cable length asymmetry, fb - beat frequency and v - velocity of light in the cable.

The linear output signal canbe obtained for  0< \theta_{b}<\pi

For our purpose in FOL, if we would like to measure beat frequency over a bandwidth of 200MHz, this would correspond to a cable length difference of 0.5 m (assuming the speed of light in the coaxial cable is ~ 2x108m/s.

ELOG V3.1.3-